The final answer should be in the format of "tolerance of ± in the third significant digit.


1.The drag racer P starts from rest at the start line S and then accelerates along the track. When it has traveled 102 m, its speed is 51 m/s. For that instant, determine the values of r˙ and θ˙ relative to axes fixed to an observer O in the grandstand G as shown.


28 meters from point O ( it is the base of the triangle)


r˙ = ___ m/s


θ˙ = ____ rad/s





1
Expert's answer
2021-04-01T18:34:45-0400

Given,

At a distance of 102m, speed = 51 m/s

Angle with the horizontal "=\theta"

Distance between O and S = 28m



Initial velocity = 0

"v=u+at"

"at= 51"

"\Rightarrow \overrightarrow{r}= -102\hat{i}+28\hat{j}"


"\Rightarrow |\overrightarrow{r}|=\sqrt{(-102)^2+(28)^2}"

"\Rightarrow|\overrightarrow{r}|=105.77"

"\tan(\theta)=\frac{28}{102}"


"\Rightarrow \theta=\tan^{-1}(\frac{28}{102})"

"\Rightarrow \theta = 15.35^\circ"